OpenStudy (anonymous):

May I have some help? I shall reward someone with a medal and become a fan. :-) The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (anonymous):

@iambatman hi batman, I need some help.

OpenStudy (anonymous):

not like the joker, but y'know.

OpenStudy (anonymous):

A = L W = 280 L - 5 = W + 1 which would make it square

OpenStudy (anonymous):

Ok, now, that's the answer for Part A: 280 = ( L - 5 ) ( W + 1 ) ?

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

280 = L - 5 W + 1 = 280?

OpenStudy (anonymous):

Mr. @douglaswinslowcooper ? I need some help please. :)

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

@zepdrix I'm so sorry for doing this to you again ;v;

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

We do not know that the altered dimensions would have the same area of 280, but we do not need that as we have two equations and two unknowns.

OpenStudy (anonymous):

So what would I put for Part A? And how would I find the width? :-)